### v2

@@ III-A3 Information feedback
$\min_{\begin{subarray}{c}\bm{\alpha},\,\bm{\beta}\end{subarray}}\,$ $\sum_{(i,j)\in\mathcal{L}}C_{ij}\alpha_{ij}+\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}b_{i}^{k,s}(\bar{\bm{x}})\beta_{i}^{k,s}$ (TR-D)
s. t. $\lambda_{s}^{k}\alpha_{ij}-\beta_{i}^{k,s}+\beta_{j}^{k,s}\geq 0,$
@@
$\alpha_{ij}\geq 0,~{}\forall~{}(i,j)\in\mathcal{L}.$
@@
meaning that point $(\bar{\bm{x}},\bar{\bm{y}})$ cannot define a feasible solution of problem (NS). Notice that the extreme ray $(\bar{\bm{\alpha}},\bar{\bm{\beta}})$ satisfying (9) can be obtained if problem (TR-D) is solved by a primal or dual simplex algorithm; see [34, Chapter 3]. To remove point $(\bar{\bm{x}},\bar{\bm{y}})$ from problem (FP), we can construct the following linear inequality
As shown in the following lemma, (10) is a valid inequality for problem (NS) in the sense that (10) holds at every feasible solution of problem (NS).
Lemma 2.
Inequality (10) is valid for problem (NS).
Proof.
Observe that
$~{}~{}~{}\geq\sum_{(i,j)\in\mathcal{L}}\bar{\alpha}_{ij}\left(\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\lambda_{s}^{k}r_{ij}^{k,s}\right)+$
$~{}~{}~{}=\sum_{(i,j)\in\mathcal{L}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\left(\lambda_{s}^{k}\bar{\alpha}_{ij}-\bar{\beta}_{i}^{k,s}+\bar{\beta}_{j}^{k,s}\right)r_{ij}^{k,s}\geq 0,$
where the first inequality follows from $\bar{\alpha}_{ij}\geq 0$, (4), and (5), and the last inequality follows from (8) and the first set of inequalities in (TR-D). Therefore, (10) is valid for problem (NS). ∎
From Lemma 2, we can add (10) into problem (FP) to obtain a tightened FP problem, which is still a relaxation of problem (NS). Moreover, by (9), the current solution $(\bar{\bm{x}},\bar{\bm{y}})$ is infeasible to this tightened FP problem. As a result, we can solve the tightened FP problem again to obtain a new solution. This process is repeated until case (i) happens. The above procedure is summarized as Algorithm 1.
@@ III-B Analysis Results and Remarks
First, at each iteration of Algorithm 1, the inequality in (10) will remove a feasible point $\bar{\bm{x}}$ from problem (FP) (if it is an infeasible point of problem (NS)). This, together with the fact that the feasible binary solutions $\bar{\bm{x}}$ of problem (FP) are finite, implies the following result.
Theorem 3.
For problem (NS), Algorithm 1 with a sufficiently large IterMax will either find one of its optimal solutions (if the problem is feasible) or declare the infeasibility (if it is infeasible).
Although the worst-case iteration complexity of the proposed decomposition algorithm grows exponentially with the number of variables, our simulation results show that it usually finds an optimal solution of problem (NS) within a small number of iterations; see Section V further ahead.
Second, each iteration of Algorithm 1 needs to solve an MBLP problem (FP) and an LP problem (TR-D). It is worthwhile highlighting that these two subproblems are much easier to solve than the original problem (NS) and this feature makes the proposed decomposition algorithm particularly suitable to solve large-scale NS problems. To be specific, although subproblem (FP) is still an MBLP problem, its dimension is $\mathcal{O}(|\mathcal{V}|\sum_{k\in\mathcal{K}}\ell_{k})$, which is significantly smaller than that of problem (NS); subproblem (TR-D) is an LP problem, which is polynomial time solvable within the complexity $\mathcal{O}\left(\left(|\mathcal{L}|\sum_{k\in\mathcal{K}}\ell_{k}\right)^{3.5}\right)$ [35, Section 6.6.1].
Third, as compared with traditional two-stage heuristic algorithms [16]-[24] that solve subproblems (FP) and (TR-D) in a one-shot fashion, our proposed decomposition algorithm iteratively solves the two subproblems enabling a useful information feedback between them (as shown in Fig. 2), which enables it to return an optimal solution of problem (NS) and thus much better solutions than those in [16]-[24].
Finally, the proposed decomposition algorithm is closely related to the well-known Benders decomposition algorithm [30, Chapter 8.3]. In particular, problems (FP) and (TR-D) correspond to the master problem and the (dual) Benders subproblem, respectively, and inequality (10) corresponds to the Benders feasibility cut. The proposed algorithm can be regarded as a custom application of the general Benders decomposition framework to solve the large-scale NS problem by observing that problem (NS) can be elegantly decomposed into FP and TR subproblems and by further judiciously exploiting the special structures of the decomposed problems to allow for an efficient information feedback between them.
@@ IV-A Connectivity based Inequalities
$M^{G}_{ij}=\left\{\begin{array}[]{ll}1,&\text{if there exists a path from node$i$ to node $j$};\\ 0,&\text{otherwise}.\end{array}\right.$
@@ IV-A1 Three forms of derived inequalities
$x^{k,s}_{v}=0,,~{}\forall~{}v\in\mathcal{V}(S^{k})\cup\mathcal{V}(D^{k}),~{}\forall~{}s\in\mathcal{F}^{k},~{}\forall~{}k\in\mathcal{K}.$ (17)
@@
$\quad\forall~{}s_{0}\in\{s+1,\ldots,\ell_{k}\},~{}\forall~{}s\in\mathcal{F}^{k},~{}\forall~{}k\in\mathcal{K}\backslash\{\ell_{k}\}.$ (18)
@@ IV-A2 Relations of the inequalities
[…]
@@
$\min_{\begin{subarray}{c}\bm{x},\,\bm{y}\end{subarray}}~{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (FP-I)
@@ IV-A3 Theoretical justification of the connectivity based inequalities
As both (17) and (19) are valid inequalities, problem (FP-I) can be seen as a relaxation of problem (NS). Therefore, if $(\bm{x},\bm{y},\bm{r})$ is a feasible solution of problem (NS), then $(\bm{x},\bm{y})$ must also be a feasible solution of problem (FP-I). To prove the other direction, let $({\bm{x}},{\bm{y}})$ be a feasible solution of problem (FP-I). By Theorem 5, Eqs. (17) and (19) imply Eqs. (14)–(16). Therefore, for the solution $({\bm{x}},{\bm{y}})$, (i) by (14), there exists a path from the source $S^{k}$ to the cloud node hosting function $f_{1}^{k}$; (ii) by (15), there exists a path from the cloud node hosting function $f^{k}_{\ell_{k}}$ to the destination $D^{k}$; and (iii) by (16), there exists a path from the cloud node hosting function $f_{s}^{k}$ to the cloud node hosting function $f_{s+1}^{k}$ for all $s\in\mathcal{F}^{k}\backslash\{\ell_{k}\}$. The above, together with the assumption that $C_{ij}=+\infty$ for all $(i,j)\in\mathcal{L}$, implies that problem (TR) has a feasible solution $\bm{r}$, and thus $({\bm{x}},{\bm{y}},\bm{r})$ is feasible solution of problem (NS). ∎
Theorem 6 justifies the effectiveness of the derived connectivity based inequalities (17) and (19) (for the proposed decomposition algorithm). Specifically, if all links’ capacities in the underlying network are infinite, then equipped with the connectivity based inequalities (17) and (19), the proposed decomposition algorithm (i.e., Algorithm 1) terminates in a single iteration. Even when some or all links’ capacities are finite, the connectivity based inequalities (17) and (19) can still accelerate the convergence of Algorithm 1, as they ensure that, for a solution $(\bm{x},\bm{y})$ returned by solving (FP-I), necessary paths exist from source nodes to the nodes hosting functions in the SFC to destination nodes for all services.
@@ IV-B1 Derived inequalities
$\qquad\quad~{}\leq\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}\backslash\{\ell_{k}\}}\lambda_{s}^{k}\left(\sum_{j:(j,v)\in\mathcal{L}}r^{k,s}_{jv}\right)$
$\qquad\quad~{}\leq\sum_{j:(j,v)\in\mathcal{L}}\left(\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\lambda_{s}^{k}r^{k,s}_{jv}\right)\leq\sum_{j:(j,v)\in\mathcal{L}}C_{jv},$
where the first and last inequalities follow from (22) and (4), respectively. By (2), if $y_{v}=0$, then $x_{v}^{k,s}=0$ for all $k\in\mathcal{K}$ and $s\in\mathcal{F}^{k}$, and thus the above inequality can be further strengthened as
@@
Inequalities (23) and (25) are nonlinear due to the nonlinear terms ${(x^{k,s+1}_{v}-x^{k,s}_{v})}^{+}$ and ${(x^{k,s}_{v}-x^{k,s+1}_{v})}^{+}$. They can be equivalently linearized by first introducing the binary variables
@@
to denote ${(x^{k,s+1}_{v}-x^{k,s}_{v})}^{+}$ and ${(x^{k,s}_{v}-x^{k,s+1}_{v})}^{+}$, respectively, and then rewriting (23) and (25) as the following constraints
@@
Adding (26)–(31) into (FP-I) yields a stronger relaxation of problem (NS):
$\min_{\begin{subarray}{c}\bm{x},\,\bm{y},\\ \bm{z},\,\bm{w}\end{subarray}}~{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (FP-II)

### v3

@@ III-A3 Information feedback
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\begin{subarray}{c}\bm{\alpha},\,\bm{\beta}\end{subarray}}}}$ $\sum_{(i,j)\in\mathcal{L}}C_{ij}\alpha_{ij}+\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}b_{i}^{k,s}(\bar{\bm{x}})\beta_{i}^{k,s}$
subject to $\lambda_{s}^{k}\alpha_{ij}-\beta_{i}^{k,s}+\beta_{j}^{k,s}\geq 0,$
@@
$\alpha_{ij}\geq 0,~{}\forall~{}(i,j)\in\mathcal{L}.$ (TR-D)
@@
meaning that point $(\bar{\bm{x}},\bar{\bm{y}})$ cannot define a feasible solution of problem (NS). Notice that the extreme ray $(\bar{\bm{\alpha}},\bar{\bm{\beta}})$ satisfying (9) can be obtained if problem (TR-D) is solved by the primal or dual simplex algorithm; see [36, Chapter 3]. To remove point $(\bar{\bm{x}},\bar{\bm{y}})$ from problem (FP), we can construct the following linear inequality
Observe that
$\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\bar{\beta}_{i}^{k,s}\left(\sum_{j:(j,i)\in\mathcal{{L}}}r_{ji}^{k,s}-\sum_{j:(i,j)\in\mathcal{{L}}}r_{ij}^{k,s}\right)$
$~{}~{}~{}=\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\sum_{i\in\mathcal{I}}\bar{\beta}_{i}^{k,s}\left(\sum_{j:(j,i)\in\mathcal{{L}}}r_{ji}^{k,s}-\sum_{j:(i,j)\in\mathcal{{L}}}r_{ij}^{k,s}\right)$
=$~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\stackrel{{\scriptstyle(a)}}{{=}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\left(\sum_{(i,j)\in\mathcal{{L}}}\bar{\beta}_{j}^{k,s}r_{ij}^{k,s}-\sum_{(i,j)\in\mathcal{{L}}}\bar{\beta}_{i}^{k,s}r_{ij}^{k,s}\right)}}$
$~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{=\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\sum_{(i,j)\in\mathcal{{L}}}\left(\bar{\beta}_{j}^{k,s}-\bar{\beta}_{i}^{k,s}\right)r_{ij}^{k,s}}},$ (11)
where (a) follows from
$\sum_{i\in\mathcal{I}}\bar{\beta}_{i}^{k,s}\sum_{j:(j,i)\in\mathcal{{L}}}r_{ji}^{k,s}=\sum_{(j,i)\in\mathcal{{L}}}\bar{\beta}_{i}^{k,s}r_{ji}^{k,s}=\sum_{(i,j)\in\mathcal{{L}}}\bar{\beta}_{j}^{k,s}r_{ij}^{k,s},$
$-\sum_{i\in\mathcal{I}}\bar{\beta}_{i}^{k,s}\sum_{j:(i,j)\in\mathcal{{L}}}r_{ij}^{k,s}=-\sum_{(i,j)\in\mathcal{{L}}}\bar{\beta}_{i}^{k,s}r_{ij}^{k,s}.$
Hence,
$~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\stackrel{{\scriptstyle(a)}}{{\geq}}}}\sum_{(i,j)\in\mathcal{L}}\bar{\alpha}_{ij}\left(\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\lambda_{s}^{k}r_{ij}^{k,s}\right)+$
=$~{}~{}~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\stackrel{{\scriptstyle(b)}}{{=}}}}\sum_{(i,j)\in\mathcal{L}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\left(\lambda_{s}^{k}\bar{\alpha}_{ij}-\bar{\beta}_{i}^{k,s}+\bar{\beta}_{j}^{k,s}\right)r_{ij}^{k,s}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\stackrel{{\scriptstyle(c)}}{{\geq}}}}0,$
where (a) follows from $\bar{\alpha}_{ij}\geq 0$, (4), and (5); (b) follows from (11); and (c) follows from (8) and the first set of inequalities in (TR-D). This shows that inequality (10) is a valid inequality for problem (NS) in the sense that (10) holds at every feasible solution of problem (NS). Therefore, we can add (10) into problem (FP) to obtain a tightened FP problem, which is still a relaxation of problem (NS). Moreover, by (9), the current solution $(\bar{\bm{x}},\bar{\bm{y}})$ is infeasible to this tightened FP problem. As a result, we can solve the tightened FP problem again to obtain a new solution. This process is repeated until case (i) happens. The above procedure is summarized as Algorithm 1.
@@ III-B Analysis Results and Remarks
First, at each iteration of Algorithm 1, the inequality in (10) will remove a feasible point $\bar{\bm{x}}$ from problem (FP) (if it is an infeasible point of problem (NS)). This, together with the fact that the feasible binary solutions $\bar{\bm{x}}$ of problem (FP) are finite, implies that, for problem (NS), Algorithm 1 with a sufficiently large IterMax will either find one of its optimal solutions (if the problem is feasible) or declare the infeasibility (if it is infeasible). Although the worst-case iteration complexity of the proposed decomposition algorithm grows exponentially with the number of variables, our simulation results show that it usually finds an optimal solution of problem (NS) within a small number of iterations; see Section V further ahead.
Second, each iteration of Algorithm 1 needs to solve an MBLP problem (FP) and an LP problem (TR-D). It is worthwhile highlighting that these two subproblems are much easier to solve than the original problem (NS) and this feature makes the proposed decomposition algorithm particularly suitable to solve large-scale NS problems. To be specific, although subproblem (FP) is still an MBLP problem, both of its numbers of variables and constraints are $\mathcal{O}(|\mathcal{V}|\sum_{k\in\mathcal{K}}\ell_{k})$. This is significantly smaller than those of problem (NS), which are
$\mathcal{O}\left((|\mathcal{V}|+|\mathcal{L}|)\sum_{k\in\mathcal{K}}\ell_{k}\right)~{}\text{and}~{}\mathcal{O}\left({\min}\left\{|\mathcal{I}|\sum_{k\in\mathcal{K}}\ell_{k},|\mathcal{L}|\right\}\right),$
respectively, especially when the number of cloud nodes is much smaller than the numbers of nodes and links in the network. Subproblem (TR-D) is an LP problem, which is polynomial time solvable within the complexity $\mathcal{O}\left(\left(|\mathcal{L}|\sum_{k\in\mathcal{K}}\ell_{k}\right)^{3.5}\right)$ [37, Chapter 6.6.1].
Finally, as compared with traditional two-stage heuristic algorithms [16]-[24] that solve subproblems (FP) and (TR-D) in a one-shot fashion, our proposed decomposition algorithm iteratively solves the two subproblems enabling a useful information feedback between them (as shown in Fig. 1), which enables it to return an optimal solution of problem (NS) and thus much better solutions than those in [16]-[24].
@@ IV-A Connectivity based Inequalities
$M^{G}_{ij}=\left\{\begin{array}[]{ll}1,&\text{if there exists a path from node$i$ to node $j$};\\[5.0pt] 0,&\text{otherwise}.\end{array}\right.$
@@ IV-A1 Three forms of derived inequalities
$x^{k,s}_{v}=0,~{}\forall~{}v\in\mathcal{V}(S^{k})\cup\mathcal{V}(D^{k}),~{}\forall~{}s\in\mathcal{F}^{k},~{}\forall~{}k\in\mathcal{K}.$ (18)
@@
$\quad\forall~{}s_{0}\in\{s+1,\ldots,\ell_{k}\},~{}\forall~{}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{s\in\mathcal{F}^{k}\backslash\{\ell_{k}\}}},~{}\forall~{}k\in\mathcal{K}.$ (19)
@@ IV-A2 Relations of the inequalities
The following two relationships are true: (i) $\mathcal{X}^{1}=\mathcal{X}^{2}=\mathcal{X}^{3}$; and (ii) $\mathcal{X}^{3}_{{\text{L}}}\subseteq\mathcal{X}^{2}_{{\text{L}}}\subseteq\mathcal{X}^{1}_{{\text{L}}}$  [^1] Note that in general, it is possible that two 0-1 sets, represented by different inequalities and binary constraints $\bm{x}\in\{0,1\}^{n}$, are equivalent but their linear relaxations are different; see [39, Chapter 1.6–1.7]..
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\begin{subarray}{c}\bm{x},\,\bm{y}\end{subarray}}}}{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (FP-I)
@@ IV-A3 Theoretical justification of the connectivity based inequalities
As both (18) and (20) are valid inequalities, problem (FP-I) can be seen as a relaxation of problem (NS). Therefore, if $(\bm{x},\bm{y},\bm{r})$ is a feasible solution of problem (NS), then $(\bm{x},\bm{y})$ must also be a feasible solution of problem (FP-I).
To prove the other direction, let $({\bm{x}},{\bm{y}})$ be a feasible solution of problem (FP-I). By Theorem 3, Eqs. (18) and (20) imply Eqs. (15)–(17). Therefore, for the solution $({\bm{x}},{\bm{y}})$, (i) by (15), there exists a path from the source $S^{k}$ to the cloud node hosting function $f_{1}^{k}$; (ii) by (16), there exists a path from the cloud node hosting function $f^{k}_{\ell_{k}}$ to the destination $D^{k}$; and (iii) by (17), there exists a path from the cloud node hosting function $f_{s}^{k}$ to the cloud node hosting function $f_{s+1}^{k}$ for all $s\in\mathcal{F}^{k}\backslash\{\ell_{k}\}$. The above, together with the assumption that $C_{ij}=+\infty$ for all $(i,j)\in\mathcal{L}$, implies that problem (TR) has a feasible solution $\bm{r}$, and thus $({\bm{x}},{\bm{y}},\bm{r})$ is feasible solution of problem (NS). ∎
Theorem 4 justifies the effectiveness of the derived connectivity based inequalities (18) and (20) for the proposed Algorithm 1. Specifically, if all links’ capacities in the underlying network are infinite, then equipped with the connectivity based inequalities (18) and (20), the proposed Algorithm 1 terminates in a single iteration. Even when some or all links’ capacities are finite, the connectivity based inequalities (18) and (20) can still accelerate the convergence of Algorithm 1, as they ensure that, for a solution $(\bm{x},\bm{y})$ returned by solving (FP-I), necessary paths exist from source nodes to the nodes hosting functions in the SFC to destination nodes for all services.
@@ IV-B1 Derived inequalities
$\qquad\quad~{}\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}\backslash\{\ell_{k}\}}\lambda_{s}^{k}\left(\sum_{j:(j,v)\in\mathcal{L}}r^{k,s}_{jv}\right)$
$\qquad\quad~{}=\sum_{j:(j,v)\in\mathcal{L}}\left(\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}\backslash\{\ell_{k}\}}\lambda_{s}^{k}r^{k,s}_{jv}\right)$
$\qquad\quad~{}\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{j:(j,v)\in\mathcal{L}}\left(\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}\cup\{0\}}\lambda_{s}^{k}r^{k,s}_{jv}\right)\stackrel{{\scriptstyle(c)}}{{\leq}}\sum_{j:(j,v)\in\mathcal{L}}C_{jv},$
where (a) follows from (23); (b) follows from $r^{k,\ell_{k}}_{jv}\geq 0$ for all $k\in\mathcal{K}$ and $j\in\mathcal{I}$ with $(j,v)\in\mathcal{L}$; and (c) follows from (4). By (2), if $y_{v}=0$, then $x_{v}^{k,s}=0$ for all $k\in\mathcal{K}$ and $s\in\mathcal{F}^{k}$, and thus the above inequality can be further strengthened as
@@
Adding inequalities (24) and (26) into (FP-I) yields a stronger relaxation of problem (NS):
${\operatorname*{minimize}_{\begin{subarray}{c}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{x},\,\bm{y}}}\end{subarray}}}{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (27)
[…]
Problem (27) is a mixed integer nonlinear programming problem due to the nonlinear terms ${(x^{k,s+1}_{v}-x^{k,s}_{v})}^{+}$ and ${(x^{k,s}_{v}-x^{k,s+1}_{v})}^{+}$ in (24) and (26). However, they can be equivalently linearized. To do so, we first note that by $x_{v}^{k,s+1},x_{v}^{k,s}\in\{0,1\}$, ${(x^{k,s+1}_{v}-x^{k,s}_{v})}^{+}$, ${(x^{k,s}_{v}-x^{k,s+1}_{v})}^{+}\in\{0,1\}$ must hold. Therefore, we can introduce binary variables
@@
to denote
$z_{v}^{k,s}={(x^{k,s+1}_{v}-x^{k,s}_{v})}^{+},\,\forall\,v\in\mathcal{V},\,\forall\,s\in\mathcal{F}^{k}\backslash\{\ell_{k}\},\,\forall\,k\in\mathcal{K},$ (30)
$w_{v}^{k,s}={(x^{k,s}_{v}-x^{k,s+1}_{v})}^{+},\,\forall\,v\in\mathcal{V},\,\forall\,s\in\mathcal{F}^{k}\backslash\{\ell_{k}\},\,\forall\,k\in\mathcal{K},$ (31)
and rewrite (24) and (26) as the following constraints
@@
As a result, problem (27) is equivalent to
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\begin{subarray}{c}\bm{x},\,\bm{y},\,\bm{z},\,\bm{w}\end{subarray}}}}{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (34)
@@
We can further linearize constraints (30) and (31) as
and obtain an equivalent MBLP formulation as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\begin{subarray}{c}\bm{x},\,\bm{y},\,\bm{z},\,\bm{w}\end{subarray}}}}{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (FP-II)
[…]
Indeed, constraints (30) and (31) imply (35) and (36), respectively. On the other hand, due to the reason that variables $\{z_{v}^{k,s}\}$ and $\{w_{v}^{k,s}\}$ do not appear in the objective function of (FP-II), there always exists an optimal solution of problem (FP-II) such that (30) and (31) hold. Therefore, problems (34) and (FP-II) are equivalent.